Gas is confined in a tank at a pressure of 1.06E+6 Pa and a temperature of 14.3
Answer #1

Initial Condition:

P1 = 1.06E+6 Pa, T1 = 14.3oC = 14.3+273 = 287.3K

Say volume is V1 and number of moles of gas is n1

Final condition:

P2 = ?, T2 = 61.6oC = 334.6K

V1 = V2 (Since tank remains the same)

n_{2} = \frac{n_{1}}{2} (Since half of the gas is removed)

On using ideal gas equations for both the conditions and diving them, we get

\frac{P_{2}V_{2}}{P_{1}V_{1}}=\frac{n_{2}RT_{2}}{n_{1}RT_{1}}

\Rightarrow \frac{P_{2}}{1.06\times 10^{6}}=\frac{334.6}{2\times287.3}

\Rightarrow P_{2} = 1.06\times 10^{6}\times \frac{334.6}{2\times287.3} = 6.17\times 10^{5} Pa}

which is the new pressure in the tank.